OpenStudy (anonymous):

is (0,3) a solution of the equation x+y=3

OpenStudy (anonymous):

@phi @jcpd910 @AnswerMyQuestions

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

We need this in slope intercept form. x + y = 3 y = -x + 3

OpenStudy (anonymous):

y = mx + b m is the slope, b is the y intercept.

OpenStudy (anonymous):

0 = x and 3 = 7

OpenStudy (anonymous):

the y intercept is the point where x is 0.

OpenStudy (anonymous):

So (0,3) is a solution.

OpenStudy (anonymous):

So you plug it in. 0 + 3 = 3. Is this a true equation?

OpenStudy (anonymous):

@jcpd910 You sounded a little complicating.

OpenStudy (anonymous):

I explained it in a way that applies to all equations like this, not just to this problem.

OpenStudy (anonymous):

Ahhh